OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = (4(4n+1)(8n+7))/(6)

OpenStudy (badhi):

when n=1, $$\frac{4(4n+1)(8n+7)}{6}=\frac{4\times 5\times 15}{6}=50\not =4.6$$